Constants| Periodic Table Find the electric field 0.50 cm from the wire. A long, thin wire carrying 5.6 nC/m runs down the center of a long, thin-walled, hollow pipe with radius 1.0 cm carrying -3.6 nC/m spread uniformly over its surface Express your answer using two significant figures. kN/C Submit ue ▼ Part B Find the electric field 1.5 cm from the wire Express your answer using two significant figures. kN/C
Answer #1

Solution)

Part a)

Let the wire is infinite

We know, E=L/(2*pi*r*e)

where:

L=linear charge density (5.6E-9 C/m)

On Solving,

E= (5.6E-9)/(2*pi*0.5E-3*8.85E-12)

E=2.014E5 N/C =2.02*10^3 kN/C (Ans)

====

Part b)

E= (5.6E-9)/(2*pi*1.5E-3*8.85E-12)

E=6.7 * 10^4  N/C =67 kN/C(Ans)

=====